• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric Voutier

    Log entry time 08:01:42 on October 4,2002

    Entry number 87303

    keyword=Shift Summary


    New Crew on Shift: Eugene Chudakov, Xiaodong Jiang, Eric Voutier
    ****************

    0:00 Left arm Q1 is still having a low cryogenic level, then we will run tonight
    Moller and full HARP + Ep energy measurement

    0:15 Ask for control access because the target motion does not work; Dave has
    been paged ans is coming.

    1:25 The target motion problem is fixed, ask for beam permit but MCC has FSD
    problems with our line.

    1:36 Hall A is back to beam permit.

    1:58 MCC puts back the hall in "control access" in order to fixe FSD problem.

    2:56 Problem is fixed, the IOC of the Moller target has been rebooted.

    3:00 Hall A back to beam permit.

    3:29 Setting up the Moller quads and dipole.
    .........

    8:00 Beam trips, MCC takes beam away for checks